The moment of inertia of a thin spherical shell of mass M and radius r, about a daimeter is Mr^2 . Its radius of gyration k about tangent will be

Grade:11

1 Answers

Agrata Singh
5 years ago
This can be solved by parallel axis theorem whisch states that
IA = ICM + Ma2
Where IA is the MI abt reqd axis ICM is the MI abt Centre of mass and a is the perpendicular distance between axis A and CM
Here IA is the MI about a tangent so a = R
IA = MR2 + MR2 = 2MR2
Now radius of gyration k is found from the formula
Mk2 = IA  =2MR2
k2 = 2R2
Hope this helps you